ROVIBRONIC BANDS OF THE Ã − X̃ ELECTRONIC TRANSITION OF CH_{3}O_{2} AND CD_{3}O_{2} DETECTED WITH CAVITY RINGDOWN ABSORPTION IN THE NIR.

\underline{PATRICK RUPPER}, ERIN N. SHARP and TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, 120 W. 18th Avenue, Columbus OH 43210.

Methyl peroxy (CH_{3}O_{2}) is the simplest alkyl peroxy radical. It is therefore the starting point for the subsequent investigation of larger alkyl peroxy radicals, which are key reaction intermediates in a variety of processes in atmospheric as well as in combustion chemistry. Its spectroscopic characterization is a prerequisite for monitoring its presence and concentration in order to understand as completely as possible the chemistry of gas phase reactions involving these reactive species. Several investigations of the spectrum of methyl peroxy have recently been reported^{a}^{,}^{b}.

We now report several rovibronic bands of CH_{3}O_{2} and CD_{3}O_{2} in their Ã^{2}A′ - X̃^{2}A′′ electronic transitions in the range 1.22 - 1.40 μm with the cavity ringdown technique that have not been discussed in detail before. While the electronic origins and COO symmetric stretching modes for these two species have been reported previously, other rovibronic bands are described here for the first time. These bands include the COO bending modes as well as many hot bands involving the methyl torsional mode (ν_{12}). Both typical sequence structure and some atypical transitions involving ν_{12} have been observed. The latter ones are due to the location of the vibrational energy levels with respect to the relatively low, but quite different, barriers in the X̃ and Ã states. The 12_{2}^{2} band in CH_{3}O_{2} and the 12_{3}^{3} band in CD_{3}O_{2} show quite different structures than the origin bands, an effect which results from tunneling splittings comparable to the rotational contour.

aM. B. Pushkarsky, S. J. Zalyubovsky and T. A. Miller J. Chem. Phys. 112, 10695 (2000).

bC.-Y. Chung, C.-W. Cheng and Y.-P. Lee The 61st Symposium on Molecular Spectroscopy, 2006, RF02.